Consider the following hypotheses: Ho: μ s 57.1 HA: μ > 57.1 A sample of 22 observations yields a sample mean of 58.2. Assume that the sample is drawn from a normal population with a population standard deviation of 3.6. (You may find it useful to reference the appropriate table: z table or t table) a-1. Find the p-value. pvalue 20.10 Op-value0.01 0.01 s p-value<0.025 0.025 s p-value 0.05 0.05 s p-value<0.10a-2. what is the conclusion if α-010? O Do not reject Ho since the p-value is smaller than a. Do not reject Ho since the p-value is greater than a o Reject Ho since the p-value is smaller than a. OReject Ho since the p-value is greater than a a-3. Interpret the results at a 0.10 We cannot conclude that the sample mean is greater than 571 We conclude that the sample mean is greater than 571. We cannot conclude that the population mean is greater than 571 We conclude that the population mean is greater than 571.b-1. Calculate the p-value if the above sample mean was based on a sample of 108 observations. O p-value< 0.01 0.01 s p-value0.025 0.025 s p-value < 0.05 0.05 s p-value < 0.10 Opvalue 2010 b-2, what is the conclusion if α-0.10? Reject Ho since the p-value is smaller than a 0 Reject Ho since the p-value is greater than a. Do not reject Ho since the p-value is smaller than a Do not reject Ho since the p-value is greater than ab-3. Interpret the results at a-0.10 We conclude that the sample mean is greater than 571 We cannot conclude that the sample mean is greater than 571 We conclude that the population mean is greater than 571 We cannot conclude that the population mean is greater than 571
Answer #1

Solution:

A sample of 22 observations yields a sample mean of 58.2. Assume that the sample is drawn from a normal population with a population standard deviation of 3.6

a-1. Find the p-value.

Answer: 0.05 〈 p-value 〈 0.10

Explanation:

To find the p-value we need to find the test statistic.

The test statistic is:

large z=rac{ar{x}-mu}{rac{sigma}{sqrt{n}}}

8.2-57.1 3.6 V22

4 1

Now, the p-value is:

p-ualue = P(z > 1.43)

Using the standard normal table, we have:

p-value = P(z > 1.43) = 0.0764

Therefore, the  0.05 〈 p-value 〈 0.10

a-2. What is the conclusion if α = 0.10 ?

Answer: Reject H0 since the p-value is smaller than alpha

a-3. Interpret the results at α = 0.10

Answer: We conclude that the population mean is greater than 57.1

b-1. calculate the p-value if the above sample mean was based on a sample of 108 observations.

Answer: p - value < 0.01

Explanation:

To find the p-value we need to find the test statistic.

The test statistic is:

large z=rac{ar{x}-mu}{rac{sigma}{sqrt{n}}}

8.2-57.1 3.6 V108

large =3.18

Now, the p-value is:

p_ ualue = P(z > 3.18)

Using the standard normal table, we have:

p-ualue P(z > 3.18) 0.0007

Therefore, the  p - value < 0.01

b-2. What is the conclusion if α = 0.10?

Answer: Reject H0 since the p-value is smaller than alpha

b-3. Interpret the results at α = 0.10

Answer: We conclude that the population mean is greater than 57.1